# VoxAct-B: Voxel-Based Acting and Stabilizing Policy for Bimanual Manipulation

### v1

@@ 6.2 Physical Results
Figure 4 shows real-world examples of VoxAct-B. In Open Drawer, success is when the stabilizing arm holds the drawer from the top while the acting arm pulls the top part. VoxAct-B succeeds in 6 out of 10 trials; the failures include robot joints hitting their limits, imprecision in grasping the handle, and collisions with the drawer. In Open Jar, a success is when the stabilizing arm grasps the jar while the acting arm unscrews the lid. VoxAct-B succeeds in 5 out of 10 trials. While the stabilizing arm performs well in grasping the jar (9 out of 10 successes), the acting arm struggles with unscrewing the lid, succeeding only 5 out of 10 times due to imprecise grasping of the lid.
@@ Appendix A Simulation Benchmark for Bimanual Manipulation
A.1 Additional Simulation and Task Details
@@
Moreover, we modify Open Jar, Open Drawer, and Put Item in Drawer to support bimanual manipulation: (1) adding an additional Franka Panda arm with a wrist camera; (2) adding new waypoints for the additional arm; (3) adjusting the front camera’s position to capture the entire workspace; (4) removing the left shoulder, right shoulder, and overhead cameras. The new tasks use a three-camera setup: front, left wrist, and right wrist. We also modify the data generation pipeline to use motion planning with the new waypoints, process RGB-D images and the new arm’s proprioception data (joint position, joint velocities, gripper open state, gripper pose), and include the $[x,y,z]$ position (world coordinates) of the object of interest.
@@
The success conditions of these tasks have also been modified: for Open Jar, we define a proximity sensor in the jar bottle to detect whether an arm has a firm grasp of the jar (the gripper’s opening amount is between 0.5 and 0.93); for Open Drawer, we define a proximity sensor on the top of the drawer to detect whether an arm is stabilizing the drawer. While a robot arm could still “open” the drawer without the other arm’s stabilization, we would not classify it a success in Open Drawer.
@@ Appendix B Real-World Experimental Details
Appendix C Additional Implementation Details
@@ C.1 Baselines
Diffusion Policy 21.3 46.7 28.0 56.0 14.7 22.7
ACT w/Transformers 30.7 57.3 32.0 37.3 36.0 82.7
Table 4: Ablation results of ACT and Diffusion Policy trained on 100 demonstrations. “FAS” refers to the demonstrations with fixed acting and stabilizing arms (i.e., right acting and left stabilizing), while “FAS+NSV” refers to fixed acting and stabilizing and without size variation in the environment. We use the same validation and test data as the baseline comparison.

### v2

@@ 6.2 Physical Results
Figure 4 shows real-world examples of VoxAct-B. In Open Drawer, success is when the stabilizing arm holds the drawer from the top while the acting arm pulls the top part. VoxAct-B succeeds in 6 out of 10 trials; the failures include robot joints hitting their limits, imprecision in grasping the handle, and collisions with the drawer. In Open Jar, a success is when the stabilizing arm grasps the jar while the acting arm unscrews the lid. For the experiment with fixed roles of each arm, VoxAct-B succeeds in 5 out of 10 trials. While the stabilizing arm performs well in grasping the jar (9 out of 10 successes), the acting arm struggles with unscrewing the lid, succeeding only 5 out of 10 times due to imprecise grasping of the lid. For the experiment with unfixed roles of each arm, we train VoxAct-B on 10 left-acting, right-stabilizing and 10 right-acting, left-stabilizing demonstrations. It succeeds in 5 out of 10 trials, demonstrating its ability to learn from multi-modal, real-world data.
@@ 7 Conclusion
Acknowledgments
We thank our colleagues Gautam Salhotra, Kr Zentner, Yigit Korkmaz, Yunshuang Li, and Guangyao Shi for helpful writing feedback.
@@ Appendix A Simulation Benchmark for Bimanual Manipulation
A.1 Additional Implementation Details
[…]
Deciding the best checkpoints for VoxAct-B and ablations is nontrivial since iterating over all possible combinations is computationally expensive. For example, with 400,000 training steps, using the same 10,000 checkpoint interval means there are $40\times 40=1600$ possible combinations. Therefore, with the validation data, we use the latest stabilizing checkpoint to evaluate all acting checkpoints; we use the best acting checkpoint to evaluate all stabilizing checkpoints. Then, we use the best-performing acting and stabilizing checkpoints to obtain the test success rate.
The policy is trained with a batch size of 1 on an Nvidia 3080 GPU for two days. Note that the batch size is not optimized based on GPU memory capacity. The batch size can be increased to 6 using an Nvidia RTX Titan with 24 GB of VRAM. Moreover, VoxAct-B is much more efficient than Bimanual PerActs, requiring only two days of training to reach one million training iterations, while it would take Bimanual PerActs ten days to reach the same number of training iterations.
During policy evaluation, the language goal, $\ell_{as}$ or $\ell_{sa}$, is determined by VLMs based on the given task. Specifically, in Open Drawer and Put Item in Drawer, we use the drawer’s pose relative to the front camera to determine which robot arm the drawer is facing. If it faces the left robot arm, $\ell_{as}$ is selected because the orientation gives the left (acting) arm a better angle for opening the drawer, and vice versa for the right robot arm. In Open Jar, we use the jar’s pose to determine which robot arm it is closer to. If it is closer to the left robot arm, $\ell_{as}$ is selected because it provides the right arm with a better angle for grasping the jar based on our experience with this task. This also applies to the block in Hand Over Item.
A.2 Additional Simulation and Task Details
@@
Moreover, we modify Open Jar, Open Drawer, and Put Item in Drawer to support bimanual manipulation: (1) adding an additional Franka Panda arm with a wrist camera; (2) adding new waypoints for the additional arm; (3) adjusting the front camera’s position to capture the entire workspace; (4) removing the left shoulder, right shoulder, and overhead cameras. The new tasks use a three-camera setup: front, left wrist, and right wrist. We also modify the data generation pipeline to use motion planning with the new waypoints, process RGB-D images and the new arm’s proprioception data (joint position, joint velocities, gripper open state, gripper pose), and include the $[x,y,z]$ position (world coordinates) of the object of interest. These modifications also apply to the Hand Over Item task.
@@
The success conditions of these tasks have also been modified: for Open Jar, we define a proximity sensor in the jar bottle to detect whether an arm has a firm grasp of the jar (the gripper’s opening amount is between 0.5 and 0.93); for Open Drawer, we define a proximity sensor on the top of the drawer to detect whether an arm is stabilizing the drawer. While a robot arm could still “open” the drawer without the other arm’s stabilization, we would not classify it a success in Open Drawer. Similarly, for Open Jar, a robot arm could “open” the jar lid without a firm grasp on the jar from the other arm, but this would not be classified as a success.
For the additional task variations, each method must learn from multi-modal demonstrations. For example, given ten training data, half have left-acting and right-stabilizing demonstrations, and the other half have right-acting and left-stabilizing demonstrations. During evaluation, the drawer from the drawer tasks is randomly rotated to face the left arm in the first half of the episodes and the right arm in the last half. This variation can make it kinematically infeasible for the left or right robot arm to open the drawer, requiring each method to determine the appropriate arm for acting and stabilizing. Additionally, we randomly scale objects (drawers and jars) from 90% to 100% of their original size.
Open Open Put Item
Jar Drawer in Drawer
ACT w/Transformers 2.7 12.0 14.7
VoxAct-B (ours) 21.3 62.7 17.3
Table 3: Multi-task results of ACT and VoxAct-B trained on 10 demonstrations of each task and evaluated across three training seeds.
A.3 Multi-Task Experiment
We train a single policy of ACT and VoxAct-B on Open Jar, Open Drawer, and Put Item in Drawer, with 10 demonstrations for each task. For evaluation, both methods use the checkpoint with the best average validation success rate across all tasks. We use the same validation and test data as the baseline comparison. Table 3 presents the multi-task results, showing that VoxAct-B outperforms ACT on all tasks.
A.4 Automatic Selection of $\alpha$
Instead of using a predefined $\alpha$, we use the VLMs to automatically determine the value of $\alpha$ by computing the largest dimension of the object of interest, with a small padding added. The method using the estimated $\alpha$ achieved a success rate of 32% on Open Drawer, compared to 73.6% when using the predefined $\alpha$, evaluated across three training seeds. We suspect the performance drop is due to the additional voxel resolutions the method must learn. As $\alpha$ varies between demonstrations due to the rescaling of objects, voxel resolution $(\frac{L}{\alpha x},\frac{W}{\alpha x},\frac{H}{\alpha x})$ also changes, making policy learning more challenging.
A.5 Jars and Drawers of Different Appearances
We introduce two new jars and drawers of different appearances. In the new Open Jar experiment (i.e., different from the Open Jar experiment in the main paper), each method is trained on 10 demonstrations using 3 different jars. The two new jars are imported from the PartNet-Mobility Dataset [72]. Each method is then evaluated on 25 episodes of validation and test data using these jars across three training seeds. ACT obtains a success rate of 0%, while our method achieves a success rate of 41.3%. In the new Open Drawer experiment, each method is trained on 10 demonstrations using a drawer with 3 different textures, as we were unable to find a drawer with similar structures online. This experiment follows the same setup as the new Open Jar experiment. In summary, ACT achieves a success rate of 9.3%, while our method achieves a success rate of 33.3%.
A.6 Ablation: Single Camera Setup
Instead of using a three-camera setup, we use a (front) single-camera setup on Open Drawer. In this setup, as the end-effector approaches the drawer handle, the view may be obstructed by the robot arm. However, with the addition of a wrist camera, the handle remains visible, allowing the voxel grid to be constructed with a clear view of the drawer handle. This ablation tests how well VoxAct-B handles occlusion with a limited number of camera views. With a single-camera setup, VoxAct-B achieves a success rate of 60%, while using the original camera setup (front, left wrist, and right wrist) yields a success rate of 73.3%, evaluated across three training seeds.
@@ Appendix B Real-World Experimental Details
Appendix C Additional Implementation Details for the Baselines
@@
Diffusion Policy 20.8 40.8 24.0 46.4 14.4 19.2
ACT w/Transformers 31.2 56.0 28.8 35.2 34.4 75.2
Table 5: Ablation results of ACT and Diffusion Policy trained on 100 demonstrations and evaluated across five training seeds. “FAS” refers to the demonstrations with fixed acting and stabilizing arms (i.e., right acting and left stabilizing), while “FAS+NSV” refers to fixed acting and stabilizing and without size variation in the environment. We use the same validation and test data as the baseline comparison.